OpenStudy (anonymous):

Let f(y) = e^y + e^(2y). Find f ', the derivative of f.

OpenStudy (anonymous):

using prodcuct rule - do you know?

OpenStudy (anonymous):

No, I don't think that's it.

OpenStudy (anonymous):

lets try this \[ f(y) = e^y + e^{2y}\] \[f'(y)=e^y+2e^{2y}\]

OpenStudy (anonymous):

first because \[\frac{d}{dx}e^x=e^x\] second by the chain rule

OpenStudy (anonymous):

satellite - so the 1st derivative of e^x is still e^x

OpenStudy (anonymous):

 why wouldd the 2nd term not be 2e^y...